At the end of the current year, the accounts receivable account has a debit balance of $895,000 and sales for the year total $10,150,000.

a. The allowance account before adjustment has a credit balance of $12,100. Bad debt expense is estimated at 1/2 of 1% of sales.

b. The allowance account before adjustment has a credit balance of $12,100.An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $38,700.

c. The allowance account before adjustment has a debit balance of $7,400. Bad debt expense is estimated at 1/4 of 1% of sales.

d. The allowance account before adjustment has a debit balance of $7,400. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $61,400.

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.
Answer #1

calculate amount of adjusting entry

a) Bad debt expense = 10150000*1%*1/2 = 50750

b) bad debt expense = 38700-12100 = 26600

c) Bad debt expense = 10150000*1%*3/4 = 76125

d) bad debt expense = 61400+7400 = 68800